OpenStudy (anonymous):

What is the best way to classify each equation 1.22-3x+7x=4(x+5) A. identity B contradiction C. neither

OpenStudy (anonymous):

@shinalcantara

OpenStudy (chosenmatt):

contradiction because the ideas and the features are opposed on another

OpenStudy (shinalcantara):

22 - 3x + 7x = 4(x+5) 22 + 4x = 4x + 20 if you try to simplify further you'll have 4x - 4x = 20-22 you'll end with 0 and -2 and there's no solution to this. B